Find f(x-h) when f(x) = x^2 +2x -7
Thank you for showing me the steps.

Answer by Edwin McCravy(20054)


f(x) = x² + 2x - 7

Spread it way apart like this:

f(  x  ) =   x  ² + 2  x   - 7

Erase all the x's

f(     ) =      ² + 2      - 7

Put (x-h) where the x's were

f((x+h)) = (x+h)² + 2(x+h) - 7

We don't need double parentheses on the left,
so we make them single parentheses:

  f(x+h) = (x+h)² + 2(x+h) - 7

You can leave it like that or you can multiply 
the expressions out on the right.

  f(x+h) = x² + 2hx + h² + 2h + 2x - 7

Edwin